There are 5 red balls and 2 green balls. What is the probability of picking a green ball if a ball is taken initially without replacing?


1
Expert's answer
2021-01-26T04:18:14-0500

Let "E_1,E_2" and "A" be three events such that : "E_1=" Initially taken ball is red

"E_2=" Initially taken ball is green

"A=" second drawn ball is green

Then "P(E_1)=\frac{5}{7},P(E_2)=\frac{2}{7}" and "P(A/E_1)=\frac{2}{6}=\frac{1}{3},P(A/E_2)=\frac{1}{6}"

Therefore the required probability is "P(A)=P(E_1).P(A\/E_1)+P(E_2).P(A\/E_2)"

"=\frac{5}{7}.\frac{1}{3}+\frac{2}{7}.\frac{1}{6}=\frac{5}{21}+\frac{1}{21}=\frac{6}{21}=\frac{2}{7}"